320 6. A defoliant is an herbicide that removes leaves from trees and growing plants. One ionic compound used for this purpose is 11% magnesium, 37.083% chlorine, and 50.206% oxygen. What is the empirical formula for this compound? What do you think its chemical name is? (Consider the possibility that this compound contains more than one polyatomic ion) 27. O830 r c 12, 211 O, 5020o X loog 50.20 logr I wasnt able to mass ecch element to each ele mert 137
Answer #1

Answer :

Assuming a mass of sample analyzed to be 100 g we have masses of each element found as,

Mg = 12.711 g, Cl = 37.083 g and O = 50.206 g.

On dividing these massess by respective atomic massess we get corresponding # of moles of each element in unknown compound.

Atomic masses are Mg (24.306 g/mole), Cl (35.455 g/mole) and O (15.999 g/mole)

# of moles of Mg = 12.711 / 24.306 = 0.523

# of moles of Cl = 37.083 / 35.455 = 1.046

# of moles of O = 50.206/15.999 = 3.138

Mole ratio hence is,

Mg : Cl : O = 0.523 : 1.046 : 3.138

On dividing ratio throughout by 0.523 (smalles number) we get,

Mg : Cl : O = 0.523 / 0.523 : 1.046 / 0.523 : 3.138 / 0.523

Mg : Cl : O = 1 : 2 : 6 (approximated)

Hence an empirical formula is : MgCl2O6.

Name : Magnesium dichloride hexaoxide.

===================XXXXXXXXXXXXXXXX==========================